OpenStudy (anonymous):

A bicycle has tires of radius 0.35 meters. If the bicycle is traveling at a constant speed of 7.8 m/s, at approximately what angular speed are the tires rotating?

OpenStudy (anonymous):

Circumference of tire?

OpenStudy (anonymous):

2Pi r , right?

OpenStudy (anonymous):

yah.

OpenStudy (anonymous):

first I use formual V= r omega to find omega. then I conver rad/s to rev/ min. But i did not get right answer :(

OpenStudy (anonymous):

\[ \left( \frac{7.8m}{s} \right) \left( \frac{2 \pi rad}{2 \pi r}\right) = \frac{7.8}{r}rad/s\] Why did you convert to rev/min?

OpenStudy (anonymous):

that sounds like you did it right

OpenStudy (anonymous):

does the answer have to be in rev/min?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I donet get why you use (2pi rad/ 2pi r)

OpenStudy (anonymous):

That's basically the formula you used; I was just doing a conversion. 2 pi radians per circumference. It's the V = r omega

OpenStudy (anonymous):

the answer is 214

OpenStudy (anonymous):

I get 213...

OpenStudy (anonymous):

if you do like that you will have omega = 22.2 rad/ s

OpenStudy (anonymous):

\[ \frac{7.8m/s}{.35m} \cdot \frac{ 1 \ rev}{2 \pi rad} \cdot \frac{60 s}{1 \ min} = 213 \]

OpenStudy (anonymous):

that is the way i did , but i did not get right answer. I still dont understand

OpenStudy (anonymous):

ok I got it

OpenStudy (anonymous):

still dunno why it's 1 off tho... :P

OpenStudy (anonymous):

no i got it now

OpenStudy (anonymous):

nice ^_^

OpenStudy (anonymous):

thanks I have some more question.

OpenStudy (anonymous):

go for it ^^

OpenStudy (anonymous):

a bicycle travels 141 m along a circular track of radius 15 m what is the angular displacement in radians of the bicycle from its strating postion ?

OpenStudy (anonymous):

and a grindstone, initially a rest is given a constant angular acceleration so that is makes 20 rev in the first 8s what is angualar acceleration? (ans : 3.93 rad/S^2)

OpenStudy (anonymous):

What's your try for the bike? Your intuition has been spot on for everything so far :)

OpenStudy (anonymous):

:) bc bicycle is harder to ride than the car haaaaa

OpenStudy (anonymous):

if you dont know it, we can skip it

OpenStudy (anonymous):

^_^ lol I know it, just thought you might have a go. Circular motion is all pretty much conversion questions. What do you get when you divide the distance traveled by the bike by the circumference of the track?

OpenStudy (anonymous):

the radius right

OpenStudy (anonymous):

?? nope. Try again.

OpenStudy (anonymous):

omg i have not idea

OpenStudy (anonymous):

 You get the number of revolutions the bike has completed